Christian Goldbach stated that every odd number greater than 7 can be written as the sum of three prime numbers. Rewrite each of the following as the sum of three prime numbers:63, 109, and 205.
Answer by Edwin McCravy(20055)

 63 = 11+23+29 
109 = 19+43+47 
205 = 59+67+79

Edwin